Find the coefficient of x^7 y^9 in the expansion of (4x+ 5y)^16.
1
Expert's answer
2020-11-17T17:01:42-0500

Expansion of "(a+b)^n" gives us "(n+1)" terms which are given by binomial expansion "\dbinom{n}{r}a^{(n-r)}b^r" , where "r" ranges from "n" to 0.

Note that powers of "a" and "b" add up to "n" and in the given problem this "n=7+9=16".

In "(4x+5y)^{16}" , we need coefficient of "x^7y^9" , we have "7^{th}" power of "x" and as such "r=16-7=9"

and as such the desired coefficient of "x^7y^9" is given by

"\dbinom{16}{9}(4x)^{(16-9)}(5y)^9=\dfrac{16!}{9!(16-9)!}(4x)^7(5y)^9=11440*16384x^7*1953125y^9= 3.6608E14x^7y^9"

So6 the coefficient is 3.6608E14